Solve the system of equations. (If the system is dependent, enter a general solution in terms of c. If there is no solution, enter NO SOLUTION.)
2x + y - z = -9
3x + 2y + 3z = 29
7x + 4y + z = 11
(x, y, z) =

Answer by TimothyLamb(4379)

2x + y - z = -9
3x + 2y + 3z = 29
7x + 4y + z = 11
---
put the system of linear equations into standard form
---
2x + y - z = -9
3x + 2y + 3z = 29
7x + 4y + z = 11
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
This linear system is singular. It has no solution
---
system matrix
2 1 -1
3 2 3
7 4 1
---
determinant of system matrix = 0
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php